50kg of sweet can be served for 200 people . how much sweet is required for 200 peaple ?

Answer by checkley79(3341)

50/200=X/200
200X=50*200
200X=10,000
X=10,000/200
X=50 ANS.
ARE YOU SURE THIS IS THE CORRECT PROBLEM STATED????